7x-6y=-8
x-9y=7
(x,y)=

Answer by haileytucki(390)

7x-6y=-8_x-9y=7
Since -9y does not contain the variable to solve for, move it to the right-hand side of the equation by adding 9y to both sides.
7x-6y=-8_x=9y+7
Replace all occurrences of x with the solution found by solving the last equation for x. In this case, the value substituted is 9y+7.
7(9y+7)-6y=-8_x=9y+7
Multiply 7 by each term inside the parentheses.
63y+49-6y=-8_x=9y+7
Since 63y and -6y are like terms, add -6y to 63y to get 57y.
57y+49=-8_x=9y+7
Since 49 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 49 from both sides.
57y=-49-8_x=9y+7
Subtract 8 from -49 to get -57.
57y=-57_x=9y+7
Divide each term in the equation by 57.
(57y)/(57)=-(57)/(57)_x=9y+7
Simplify the left-hand side of the equation by canceling the common factors.
y=-(57)/(57)_x=9y+7
Simplify the right-hand side of the equation by simplifying each term.
y=-1_x=9y+7
Replace all occurrences of y with the solution found by solving the last equation for y. In this case, the value substituted is -1.
y=-1_x=9(-1)+7
Multiply 9 by each term inside the parentheses.
y=-1_x=-9+7
Add 7 to -9 to get -2.
y=-1_x=-2
This is the solution to the system of equations.
y=-1_x=-2